### v1

# Integrated soliton microcombs beyond the turnkey limit

Abstract
Soliton microcombs generated in optical microresonators are accelerating the transition of optical frequency combs from laboratory instruments to industrial platforms. Self-injection locking (SIL) enables direct driving of soliton microcombs by integrated lasers, providing turnkey initiation and improved coherence, but it also pins the pump close to resonance, limiting both spectral span and tuning flexibility. Here we theoretically and experimentally demonstrate that introducing a thermally tunable auxiliary microresonator extends the bandwidth of SIL soliton microcombs. By engineering hybridization of the pumped resonance, we achieve deterministic access to single-soliton states and then push operation into a far-detuned regime inaccessible to direct initiation. The resulting combs reach a near-200-nm span at a 25-GHz repetition rate, while preserving the SIL-enabled noise suppression throughout. Moreover, the added degree of freedom afforded by the coupled-resonator architecture enables orthogonal control of the comb’s repetition rate and center frequency. These advances expand the spectral reach and controllability of integrated soliton microcombs for information processing and precision metrology.
preprint: APS/123-QED
Interest in deploying optical frequency combs in practical settings—spanning metrology, communications, and computing—is rapidly growing 1. Kerr microcombs, generated in high-$Q$ optical microresonators via Kerr-driven four-wave mixing, represent a leading route to this goal 2. A particularly important class of Kerr microcombs is the soliton microcomb, in which dissipative Kerr solitons yield femtosecond pulses and phase-locked, equally spaced comb lines with a well-defined spectral envelope 3, 4, 5, 6, 7. Such sources can provide bidirectional optical–microwave links for optical clocks 8, pulse trains for precision ranging 9, 10, 11, and multiwavelength carriers for high-speed data transmission 12, 13. While initial demonstrations of soliton microcombs employed simple single-ring microresonators, their performance, particularly in terms of power efficiency and spectral span, has been substantially enhanced by incorporating more advanced resonator geometries and coupling schemes 14, 15, 16, 17, 18, 19.
Fig. 1: Dynamics of self-injection locking (SIL) in coupled microresonators. a, Schematic of a coupled-resonator device directly pumped by a distributed-feedback (DFB) laser. The original NR mode is indicated by dashed lines, and the NR (red) and AR (blue) fractions of the hybridized mode are encoded in color. b, Effect of pump-mode hybridization on the NR integrated dispersion (middle) and on the parametric gain spectrum (bottom). Both quantities are normalized to half the NR linewidth, $\kappa_{\mathrm{NR}}/2$. c, Phase diagram of soliton existence under SIL conditions, plotted as a function of comb detuning and mode shift. Arrows indicate the pathway from turnkey excitation to spectrally broadened soliton microcombs.
Photonic integration has recently transformed soliton microcombs: by directly coupling on-chip lasers or gain chips to microresonators, soliton states can form autonomously via self-injection locking (SIL) 20, 21, 22, 23, 24, 25, 26, 18, 27. Beyond turnkey operation, SIL narrows the pump-laser linewidth and allows the use of inexpensive, low-coherence lasers, enabling compact, mass-producible comb sources suitable for deployment outside the laboratory 27. However, SIL also introduces two key constraints: (i) the locking mechanism limits the achievable pump–resonator detuning, even though large detuning is essential for broadband spectra; and (ii) locking removes an independent tuning degree of freedom, whereas orthogonal control of the repetition rate and comb center frequency requires two. Photonic-crystal microresonators can partially alleviate these limitations by shifting the pump mode off the native dispersion curve, thereby permitting larger detuning while maintaining lock. In practice, however, turnkey comb initiation becomes increasingly difficult as the pump-mode shift grows, ultimately limiting the attainable detuning 18.
Here, we overcome this limitation by introducing a reconfigurable coupled-resonators architecture for integrated soliton microcombs. Although similar structures have previously been used to enhance soliton efficiency 17, this is, to our knowledge, their first application in the SIL regime. By dynamically adjusting the pump-mode frequency via an auxiliary resonator, an initiation–tuning protocol grants access to detuning values far beyond the turnkey SIL range, where markedly broader combs emerge. We experimentally implement this approach using a distributed-feedback (DFB) laser pumping Si3N4 coupled microresonators and achieve a near-200-nm span at a 25 GHz repetition rate, while preserving SIL-enabled noise suppression throughout. Moreover, the auxiliary resonator provides an additional control degree of freedom that we exploit to realize orthogonal tuning of the repetition rate and pump frequency of the comb.
Fig. 2: Generation of soliton microcombs. a, Device photographs. From left to right: Si3N4 photonic chip, DFB laser directly butt-coupled to the chip, and the fully packaged assembly in a butterfly package. b, Upper panel: measured transmission spectra at different hybridization levels, obtained by varying the AR heater power. The AR resonance, NR resonance, and hybridized resonances (AR $\times$ NR) are indicated. Lower panel: comb power versus DFB drive current over 100 consecutive scans.
Fig. 3: Tuning and characterization of soliton microcombs. a, Experimental setup. PD: photodetector; ESA: electrical spectrum analyzer; OSA: optical spectrum analyzer; FBG: fiber Bragg grating; PNA: phase-noise analyzer; LO: local oscillator; PC: polarization controller; AOFS: acousto-optic frequency shifters. b, Measured comb electrical beat note (upper panel) and DFB laser frequency (lower panel) as a function of AR heater power. c, Optical spectra of soliton microcombs in the turnkey (left) and extended (right) regimes at the operating points indicated in b. The 3-dB spectral bandwidth is extracted by fitting the spectra with $\mathrm{sech}^{2}$ envelopes. Insets: corresponding electrical beat notes of the combs. d, Power spectral density (PSD) of the frequency noise of the free-running and SIL DFB lasers. The noise floor at high offset frequencies is indicated. e, Measured 3-dB spectral-envelope bandwidth (left axis) and Lorentzian linewidth (right axis) of the microcomb as a function of AR heater power.
Fig. 4: Orthogonal control of comb properties. a, Repetition rate (red) and pump frequency (blue) via AR tuning. b, Repetition rate (red) and pump frequency (blue) via NR tuning. c, Repetition rate (red) and pump frequency (blue) via combined AR and NR tuning.
Results
Principle
Figure 1a illustrates the architecture, in which a nonlinear resonator (NR) is coupled to a bus waveguide and side-coupled to an auxiliary resonator (AR). A DFB laser is directly coupled to the bus waveguide without an optical isolator. Fabrication-induced backscattering provides optical feedback that drives SIL 28, 29, 30, 31, 32, 26, 33, and the feedback phase—the phase of the reinjected field with respect to the emitted pump—is tuned to optimize locking. At elevated pump powers, Kerr nonlinearity lifts the degeneracy between the forward- and backward-propagating modes, shifting the optimal SIL point slightly to the red of the forward resonance. This red detuning satisfies the soliton existence condition and enables turnkey comb initiation 20, 21, 24, 25, 26, but the resulting nonlinear splitting is small, limiting the accessible detuning and comb span.
[…]
Comb initiation occurs when the parametric gain exceeds the resonator loss. Increasing $\Delta$ pulls the parametric gain lobes toward $\mu=0$, merges them, and eventually suppresses the peak gain below threshold (Fig. 1b). To quantify the resulting behavior, we construct a phase diagram in the $(\zeta,\Delta)$ plane at a pump power five times the parametric-oscillation threshold of the standalone NR, subject to the SIL constraints (see Methods). We identify soliton states by scanning the feedback phase and the laser free-running frequency. Without hybridization ($\Delta=0$), the accessible comb-detuning range is narrow, with a maximum normalized detuning of 3.2 (in units of $\kappa_{\mathrm{NR}}/2$). Increasing $\Delta$ broadens the soliton region and extends the attainable detuning to 12.7 at $\Delta=8.8$. For $\Delta<8.8$, the parametric gain still exceeds the resonator loss, so solitons form automatically when the system is tuned to a preset operating point; we refer to this as the turnkey regime 24. Beyond this point, the system enters an extended regime in which direct initiation is suppressed because the parametric gain falls below the resonator loss, yet solitons seeded in the turnkey regime persist up to $\Delta=19.7$, with detuning as large as 20.6 (see Supplementary Information). Access to these larger detunings substantially broadens the comb span and relies on a continuously tunable mode shift, in contrast to photonic-crystal ring microresonators, where $\Delta$ is fixed 18.
Comb generation in the turnkey regime
[…]
Comb generation across three hybridization regimes is experimentally verified in Fig. 2b. For each heater setting, we record the transmission and comb power as functions of the DFB current, which, in the absence of SIL, sets the free-running laser frequency. Without hybridization, scanning the laser across the NR resonance predominantly yields a two-soliton state, because the maximum parametric gain occurs at multiple free-spectral-range spacings from the pump (see Supplementary Information). Nevertheless, turnkey comb generation remains possible by directly biasing the laser current to the corresponding operating point (for example, 340.5 mA). Next, the AR is tuned closer to the NR to realize partial hybridization, such that one hybridized resonance is red-shifted relative to the original NR resonance. In this condition, ramping up the laser current most frequently produces single-soliton states, providing a favorable operating point for deterministic turnkey initiation. This behavior arises because the maximum parametric gain is shifted closer to the pump and the first generated sidebands appear at mode numbers $\mu=\pm 1$ (Fig. 1b). When the AR is tuned even closer to the NR, the system approaches full hybridization, and the transmission spectrum exhibits two hybridized resonances with comparable depth. Pumping the most red-shifted hybrid mode then results in no comb formation, indicating that the system has entered the extended regime.
Comb broadening in the extended regime
As outlined in Fig. 3a, we comprehensively characterize the SIL soliton state: the optical spectrum is measured with an optical spectrum analyzer, the repetition rate is obtained from the photocurrent of a fast photodetector on an electrical spectrum analyzer, the pump frequency is tracked by heterodyning the filtered DFB output with a fixed probe laser, and the DFB frequency noise is measured using a delayed self-heterodyne interferometer and a phase-noise analyzer. The AR heater power $P_{\mathrm{AR}}$ parametrizes the AR–NR hybridization and sets the NR mode shift. A single soliton is initiated in the turnkey regime at $P_{\mathrm{AR}}\approx 120.5$ mW and subsequently driven into the extended regime by ramping up $P_{\mathrm{AR}}$. The resulting mode shift causes the DFB frequency to decrease approximately linearly due to SIL (Fig. 3b), with a gradual roll-off near the end of the tuning range indicating the approach to the maximal mode shift. In both the turnkey and extended regimes, the SIL microcomb remains a stable single soliton, as evidenced by a single, narrow repetition-rate line at 24.961 GHz. Meanwhile, the soliton spectrum broadens markedly: the 3-dB envelope bandwidth increases from 17.1 nm (turnkey) to 26.8 nm (extended), and the nonlinear conversion efficiency increases from 4.4% to 7.9% (Fig. 3c). Spectral protrusions relative to the $\mathrm{sech}^{2}$ envelope mostly arise from periodic mode hybridization induced by the interleaved AR and NR free spectral ranges, which perturb the integrated dispersion of NR. In the extended regime, the larger mode shift and increased soliton detuning reduce the impact of this distortion and smooth the central spectrum. Further engineering of the AR free spectral range to suppress periodic hybridization is expected to yield a spectrum that more closely follows a pure $\mathrm{sech}^{2}$ envelope.
A key advantage of SIL is the reduction of pump frequency noise. The achievable noise suppression depends on the microresonator quality factor and the strength of backscattering, which typically decreases at large detuning as intracavity power is reduced 26. In the coupled-resonators architecture, however, laser injection into the hybridized resonance remains efficient even at large comb detuning, so the intracavity power and associated backscattering remain sufficient to sustain noise reduction. We therefore measure the pump frequency noise in both the turnkey and extended regimes (Fig. 3d), comparing their spectra at high offset frequencies ($f\gtrsim 10~\mathrm{kHz}$) to suppress the influence of environmental and thermal noise. Within the extended regime, SIL-enabled noise suppression remains largely intact even at the maximal mode shift, with only a slight increase in the DFB frequency-noise power spectral density relative to the turnkey regime (Fig. 3d). This small increase in SIL-locked laser noise is attributable to enhanced AR–NR hybridization: with $Q_{\mathrm{NR}}\sim 4\times 10^{6}$ and $Q_{\mathrm{AR}}\sim 1\times 10^{6}$, stronger coupling lowers the $Q$ of the predominantly NR-like hybrid mode and slightly weakens linewidth compression. Figure 3e summarizes the comb properties over the entire tuning range: the 3-dB spectral-envelope bandwidth increases monotonically with $P_{\mathrm{AR}}$, while the Lorentzian linewidth, inferred from the high-frequency noise spectrum (see Methods), remains strongly suppressed relative to the free-running DFB, maintaining $20~\mathrm{dB}$ reduction.
Orthogonal control of comb properties
The coupled-microresonator architecture enables two-degree-of-freedom control of the soliton repetition rate and pump frequency by exploiting their distinct sensitivities to thermal tuning of the AR and NR. Tuning the AR modifies both the mode shift, and hence the pump frequency, and the comb detuning, thereby changing the repetition rate via Raman self-frequency shifts 34, 35, 36 and dispersive-wave recoil 5, 37. In contrast, tuning the NR predominantly shifts the pump resonance and thus the pump frequency, while only weakly perturbing the comb detuning and repetition rate; moreover, the opposing contributions from the slight change in comb detuning and the change in the FSR of the NR partially cancel. By forming appropriate linear combinations of AR and NR heater powers, these two control channels can be recombined into orthogonal eigenmodes that independently address either the repetition rate or the pump frequency.
Experimentally, we operate in the extended detuning regime with AR and NR heater powers of 127 mW and 1 mW, respectively. Tuning the AR yields an approximately linear decrease in the repetition rate of 63.9-63.9 kHz/mW and a concomitant increase in the pump frequency of 6.1 MHz/mW (Fig. 4a). Near this operating point, tuning only the NR produces a tenfold smaller change in repetition rate (6-6 kHz/mW) but a much larger pump-frequency shift of 26.2 MHz/mW (Fig. 4b). Based on these response coefficients, we implement orthogonal control by tuning the AR and NR in opposite directions with an appropriate $\sim 5{:}1$ ratio of heater-power changes (Fig. 4c). This strategy yields an effective repetition-rate tuning coefficient of 47.3-47.3 kHz/mW (normalized to $|\Delta P_{\mathrm{AR}}|+|\Delta P_{\mathrm{NR}}|$), while suppressing the pump-frequency tuning to only 0.35 MHz/mW. Although the orthogonality can be further refined to reduce residual cross-talk, the demonstrated control is already sufficient to enable potential full stabilization of the soliton microcomb 38, 39.
Discussion
In conclusion, we have introduced an integrated soliton microcomb architecture that operates beyond the conventional turnkey limit while preserving the simplicity and coherence benefits of SIL. Although demonstrated here in Si3N4 microresonators, the concept is generic and can be extended to other Kerr-comb platforms, including LiNbO3 7, 40, 41, LiTaO3 42, AlN 6, and SiON 43, among others. Several of these platforms offer intrinsic electro-optic or piezoelectric tuning, enabling fast, low-power, and programmable comb control. Even in platforms without native electro-optic functionality, heterogeneously integrated piezoelectric actuators can substantially increase the tuning speed relative to thermal heaters 44, 39. These capabilities are expected to further enhance locking performance in applications such as optical frequency division 45, 46, 47, 48, 49, optical clocks 8, and optical frequency synthesis 50.
The increased spectral span and high level of integration make this platform directly relevant to a broad range of applications. In optical frequency division, the larger comb bandwidth supports higher division ratios and therefore lower phase noise in the generated microwaves 45, 46, 47, 48, 49. In dual-comb spectroscopy, the broader spectral coverage enables simultaneous interrogation of more molecular species 51, 52, 53. If the NR free spectral range is increased to the few-hundred-gigahertz regime, the resulting comb can provide line spacings compatible with current wavelength-division-multiplexing formats for high-capacity telecommunications 12, 13. Moreover, the NR dispersion and $Q$ factor can be engineered to realize either broadband or high-power combs, tailored to specific system requirements. Overall, the presented architecture represents a significant advance over conventional SIL soliton microcombs and is well positioned for deployment across a wide range of next-generation coherent photonic systems.
Methods
Phase diagram. Reconfigurable pump-mode hybridization induced by the AR is modeled as an effective frequency shift of the pump mode within the NR mode family 17. To obtain the phase diagram in Fig. 1c, we model the SIL system with pump-mode hybridization in terms of the soliton field $\psi_{\mathrm{S}}$, the backscattered field $\psi_{\mathrm{B}}$, and the dynamics of the SIL pump laser. Their evolution is governed by the following equations of motion 54, 30, 16:
$\frac{\partial\psi_{\mathrm{S}}}{\partial\tau}$ $=-(1+i\zeta)\psi_{\mathrm{S}}+id_{2}\frac{\partial^{2}\psi_{\mathrm{S}}}{\partial\phi^{2}}+i(|\psi_{\mathrm{S}}|^{2}+2|\psi_{\mathrm{B}}|^{2})\psi_{\mathrm{S}}+i\beta\psi_{\mathrm{B}}$ (1)
$\hphantom{{}=}+i\Delta\overline{\psi_{\mathrm{S}}}+f,$
$\frac{\partial\psi_{\mathrm{B}}}{\partial\tau}$ $=-(1+i\xi)\psi_{\mathrm{B}}+i(|\psi_{\mathrm{B}}|^{2}+2P)\psi_{\mathrm{B}}+i\beta\overline{\psi_{\mathrm{S}}},$
$\xi$ $=\Delta_{\mathrm{L}}+K_{0}\mathrm{Im}\left[(1-i\alpha_{\mathrm{g}})e^{i2\phi_{\mathrm{B}}}\frac{\psi_{\mathrm{B}}}{f}\right].$
Here, $\tau$ is the normalized slow time and $\phi$ is the angular coordinate in the co-rotating frame; $\zeta$, $\xi$, and $\Delta_{\mathrm{L}}$ are the normalized comb, pump, and laser detunings (in units of $\kappa_{\mathrm{NR}}/2$), and the normalized mode shift $\Delta$ links pump and comb detunings via $\zeta=\xi+\Delta$. The parameter $d_{2}$ denotes the normalized group-velocity dispersion of the NR; $\beta$ is the dimensionless backscattering coefficient (in units of $\kappa_{\mathrm{NR}}/2$); and $|f|^{2}$ is the normalized pump power, expressed in units of the parametric-oscillation threshold of the standalone NR. $K_{0}$ is the locking bandwidth, $\alpha_{\mathrm{g}}$ the linewidth–enhancement factor, and $\phi_{B}$ the feedback phase (i.e., the phase delay between the NR and the laser). We define $P=\overline{|\psi_{\mathrm{S}}|^{2}}$ as the normalized total intracavity power and $\overline{\psi_{\mathrm{S}}}$ as the normalized amplitude of the NR pump mode. For a given $\Delta$, we sweep $\Delta_{\mathrm{L}}$ and $\phi_{\mathrm{B}}$ to determine the range of comb detunings that support solitons, holding all other parameters fixed at $|f|^{2}=5$, $\beta=0.1$, $K_{0}=50$, $\alpha_{\mathrm{g}}=0$, and $d_{2}=0.02$. To identify solitons in the turnkey regime, quantum noise corresponding to one-half photon per mode is injected into $\psi_{\mathrm{S}}$; a turnkey soliton obtained in this way is then used as the initial condition to map soliton persistence in the extended regime. A detailed derivation and simulation procedure are provided in the Supplementary Information.
Device fabrication. The Si3N4 coupled microresonators are realized on a 4-inch wafer using a subtractive fabrication process 27. First, a 786 nm-thick Si3N4 layer is deposited in two steps onto a wet-oxidized silicon substrate that incorporates stress-relief structures. The waveguide layout is then defined by electron-beam lithography, and the pattern is transferred into the Si3N4 film by dry etching. To eliminate residual N–H and Si–H bonds, the wafer is annealed at 1200C. A SiO2 cladding layer is subsequently deposited and densified in a second annealing step. Metal microheaters are patterned using a lift-off process. Finally, the wafer is diced into $5~\mathrm{mm}\times 5~\mathrm{mm}$ chips.
Laser linewidth measurement. The pump-laser frequency noise is measured with a delayed self-heterodyne interferometer. In this scheme, the laser output is divided into two paths: one path is frequency shifted by 55 MHz using an acousto-optic modulator, while the other is sent through a 1-km optical-fiber delay line. The two paths are then recombined, detected on a photodiode, and the resulting beat signal is analyzed with a phase-noise analyzer to obtain the single-sideband phase-noise spectrum $S_{\mathrm{\phi}}(f)$. The corresponding frequency-noise PSD $S_{\mathrm{F}}(f)$ is obtained via
$S_{\mathrm{F}}(f)$ $=S_{\mathrm{\phi}}(f)\frac{f^{2}}{4\sin^{2}{(\pi f\tau_{d})}},$ (2)
where $\tau_{d}$ denotes the interferometric delay time (4.8821 µs in this work). Because the denominator vanishes at offset frequencies satisfying $\sin{(\pi f\tau_{d})}=0$, the analysis in Fig. 3e is restricted to offset frequencies $(N+1/2)/\tau_{d}$, with $N$ a positive integer. The Lorentzian linewidth reported in Fig. 3e is then extracted from the high-frequency plateau of $S_{\mathrm{F}}(f)$ as
$\delta\nu$ $=2\pi\sum_{N=20}^{59}\frac{S_{\mathrm{F}}\left(\frac{N+1/2}{\tau_{d}}\right)}{30},$ (3)
and the error bars in Fig. 3e correspond to the standard deviation of $S_{\mathrm{F}}(f)$ over the discrete offset frequencies used in Eq. (3).
Feedback phase optimization. When the AR is thermally tuned in the extended regime, the feedback phase is perturbed by thermal crosstalk, which can degrade the noise-reduction performance of self-injection locking. For the Lorentzian linewidth data in Fig. 3e, we therefore re-optimize the voltage applied to the phase heater ($E_{\mathrm{phase}}$) after each AR adjustment to restore the optimal feedback phase. Specifically, we perform a fine sweep of $E_{\mathrm{phase}}$ over a narrow range with the smallest available step size, compute the corresponding Lorentzian linewidth at each setting, and choose the minimum as the operating point.
Acknowledgments
The authors thank Yinke Cheng, Xin Zhou, Jincheng Li, Zhigang Hu for assistance in fabrication, and Du Qian for assistance on sample photography. This work was supported by National Key R&D Plan of China (Grant No. 2023YFB2806702), National Natural Science Foundation of China (12293050), and the High-performance Computing Platform of Peking University. The fabrication in this work was supported by the Peking University Nano-Optoelectronic Fabrication Center, Micro/nano Fabrication Laboratory of Synergetic Extreme Condition User Facility (SECUF), Songshan Lake Materials Laboratory, and the Advanced Photonics Integrated Center of Peking University.
Author contributions
Experiments were conceived and designed by Z.W., T.X., and Q.-F.Y. Measurements and data analysis were performed by Z.W., T.X., with assistance from K.Z., X.L and B.N. Numerical simulations were performed by T.X. with assistance from Z.W. The device was designed by K.Z. and T.X. with assistance from X.L. The device was fabricated and tested by Y.W., with assistance from H.L., J.W., W.Y. and B.-B.L. The project was supervised by Q.G., Y.-F.X. and Q.-F.Y. All authors participated in preparing the manuscript.
Competing interests
Q.-F.Y., B.N., T.X., Z.W., Y.-F.X. and Q.G. are co-inventors on Chinese patent application CN202510557940.9 concerning the soliton microcomb module.
Data availability
The data that support the plots within this paper and other findings of this study are available upon publication.
Code availability
The codes that support the findings of this study are available upon publication.
Additional information
Correspondence and requests for materials should be addressed to Q-F.Y.

### v2

@@
 Experimentally, we operate in the extended detuning regime with AR and NR heater powers of 127 mW and 1 mW, respectively.
-Tuning the AR yields an approximately linear decrease in the repetition rate of 63.9-63.9 kHz/mW and a concomitant increase in the pump frequency of 6.1 MHz/mW (Fig.
+Tuning the AR yields an approximately linear decrease in the repetition rate of 63.9-63.9 kHz/mW and a concomitant increase in the pump frequency of 60.8 MHz/mW (Fig.
 4a).
-Near this operating point, tuning only the NR produces a tenfold smaller change in repetition rate (6-6 kHz/mW) but a much larger pump-frequency shift of 26.2 MHz/mW (Fig.
+Near this operating point, tuning only the NR produces a tenfold smaller change in repetition rate (6-6 kHz/mW) but a much larger pump-frequency shift of 261.7 MHz/mW (Fig.
 4b).
@@
 4c).
-This strategy yields an effective repetition-rate tuning coefficient of 47.3-47.3 kHz/mW (normalized to $|\Delta P_{\mathrm{AR}}|+|\Delta P_{\mathrm{NR}}|$), while suppressing the pump-frequency tuning to only 0.35 MHz/mW.
+This strategy yields an effective repetition-rate tuning coefficient of 47.3-47.3 kHz/mW (normalized to $|\Delta P_{\mathrm{AR}}|+|\Delta P_{\mathrm{NR}}|$), while suppressing the pump-frequency tuning to only 3.5 MHz/mW.
 Although the orthogonality can be further refined to reduce residual cross-talk, the demonstrated control is already sufficient to enable potential full stabilization of the soliton microcomb 38, 39.